• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Xiaofeng

    Log entry time 00:22:08 on November 2,2008

    Entry number 246051

    keyword=Swing shift summary for 11/01/2008

    16:00 Shift start with continue data taking on H2 target , pressure at 98 psig. Bigbite wire chambers trips a lot;
    16:50 increase beam current from 1 uA to 2 uA;
    17:20 finish data taking on H2 target, 4.9 M events
    17:35 bigbite magnet current set to 0 A, raster off
    moved to optics target position, beam current: 0.5 uA;
    18:40 Mcc took around 30 minutes for beam position tuning, and we asked for 1 uA. The CODA is not stable, roc 9 tripped.
    20:40 finish data taking for otpics study (with Bigbite sieve slit on)
    21:00 asking for controlled acess to removed bigbite sieve plate
    22:00 take 10 miutes' optics study
    22:35 taking data on H2 gas target , 1 uA beam, raster on